# Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure

### v1

@@
Abstract
@@
Dark matter-dominated cores have long been claimed for the well-studied local group dwarf galaxies. More recently, extended stellar halos have been uncovered around several of these dwarfs through deeper imaging and spectroscopy. Such core-halo structures are not a feature of conventional cold dark matter (CDM), based on collisionless particles where smooth, scale-free profiles are predicted. In contrast, smooth and prominent dark matter cores are predicted for Warm and Fuzzy/Wave Dark Matter (WDM/DM) respectively. The question arises to what extent the visible stellar profiles should reflect this dark matter core structure. Here we compare cosmological hydrodynamical simulations of CDM, WDM DM, aiming to predict the stellar profiles for these three DM scenarios. We show that cores surrounded by extended halos are distinguishable for WDM and DM, with the most prominent cores in the case of DM, where the stellar density is enhanced in the core due to the presence of the relatively dense soliton. Our analysis demonstrates that such behavior does not appear in CDM, implying that the small-scale cut-off in the power spectrum present for WDM and DM provides a core-halo transition. Consequently, we estimate the mass of the DM particle at this core-halo transition point. Furthermore, we observe the anticipated asymmetry for DM due to the soliton’s random walk, a distinctive characteristic not found in the symmetric distributions of stars in Warm and CDM models.
keywords:
Dark Matter, Galaxies: dwarfs, Galaxies: kinematics and dynamics1 Introduction
@@
Dark Matter is understood to be non-relativistic, i.e. cold, as required to form galaxies gravitationally and for explaining the spectrum of CMB fluctuations in detail (Collaboration et al., 2018). However, the standard heavy particle interpretation for Cold Dark Matter (CDM) faces a stringent laboratory absence of any new particle signature (Aprile et al., 2018; Collaboration, 2020) and several inconsistencies have emerged between CDM predictions and the properties of dwarf galaxies (Moore, 1994; de Blok, 2010; Marsh & Pop, 2015; Klypin et al., 1999; Safarzadeh & Loeb, 2021). Warm Dark Matter (WDM) is defined to be relativistic initially with a keV-scale mass that is sufficiently redshifted to become non-relativistic before galaxy formation, generating smooth cores and suppressing formation of low-mass galaxies. Alternatively, for very light bosons of , an inherently non-relativistic possibility of dark matter as a Bose-Einstein condensation (Widrow & Kaiser, 1993; Hu et al., 2000), originally termed Fuzzy Dark Matter (FDM), has been emerged by the first simulations in this context, which predict prominent solitonic standing wave cores surrounded by pervasive interference on the de Broglie wavelength scale within all galaxies and filaments (Schive et al., 2014a; Mocz et al., 2017; Veltmaat et al., 2018; Hui, 2020). This was then termed as ‘Wave Dark Matter’, DM, (Schive et al., 2014a; Hui et al., 2021). While DM predicts almost identical large-scale structure to CDM in terms of the network of filaments and clusters (Schive et al., 2014a; Mocz et al., 2017), the small-scale structure of DM halos is very different in DM paradigm, as DM cannot be confined to less than the de Broglie scale. Consequently, in DM, with a light boson mass of , the dwarf galaxy formation is suppressed and there is also a prominent solitonic core (Schive et al., 2014a, b; Mocz et al., 2017; Veltmaat et al., 2018; Niemeyer, 2020) in the galaxy.
@@
Many independent claims of dark matter and stellar cores in deep images and spectroscopy are now available for the well-studied, Local Group dwarf spheroidal galaxies, dSph, providing one of the main sources of tension for CDM for which the lack of any predicted cores reflects the scale-free nature of simplest, collisionless DM scenario. More recently the presence of extended stellar halos has been uncovered in several local dwarfs, independently, with deep imaging and spectroscopy (Chiti et al., 2021; Collins et al., 2021). A joint analysis of the available data now suggests that such a core-halo structure with a sharp density transition between the core and the halo is a common feature of both dSph galaxies and also the smaller new class of UFG (Ultra faint galaxies) (Pozo et al., 2020, 2023). This core-halo behaviour motivates the examination of hydrodynamical simulations with a comparison of the differing DM contenders as a possible distinguishing feature.
@@
For cold dark matter (CDM), the scale-free formation of DM structures is predicted to extend to arbitrarily small scales, reaching lower halo masses with higher DM concentrations. In contrast, small-scale suppression of the power spectrum is inherent to both DM and WDM. In the case of DM, this suppression limits the minimum scale of structure to below the de Broglie length, which is determined by the boson mass. Consequently, DM structures below are suppressed for a choice of boson masses of eV, motivated by the observed kpc scale of dwarf galaxy cores. This limitation has a direct consequence for the formation timescales between CDM and DM, resulting in a delay in galaxy formation for DM compared to CDM. In both cases, we observe that the formation of the first "filamentary" structures occurs earlier in CDM, where they are comprised of low-mass subhaloes. In contrast, in DM/WDM, there is no fragmentation along the filaments due to the small-scale cutoff of the power spectrum. Instead, dark matter is distributed more continuously along these filaments.
@@
The main difference between warm dark matter (WDM) and DM is that WDM exhibits a clear caustic structure in the distribution of dark matter, while DM is characterized by wave interference patterns that manifest as fringes in simulations. These fringes are now understood to generate turbulent density structures within halos, resulting from the superposition of numerous plane waves encoding the velocity dispersion in the halo (Mocz et al., 2020). In DM, these filaments persist for longer periods, as the gravitational attraction from the early formation of the first halos in cold dark matter (CDM) tends to disrupt them earlier. Moreover, the prolonged lifetimes of these filaments in wave dark matter favor the formation of a greater number of stars within them, leading to significant differences in the location and extent of stellar profiles in CDM and DM galaxies. This also causes baryonic objects to appear more diffuse or smoothed compared to CDM (Mocz et al., 2019, 2020), indicating that they could serve as excellent tracers of dark matter in DM. On the scale of solitons, the quantum pressure in DM can become sufficiently strong to counteract the self-gravity of the dark matter. This results in the formation of prominent solitonic standing waves at the base of each virialized potential, characterized by a spherical soliton core of a few kiloparsecs at the center of every dark matter halo (Schive et al., 2014a), in contrast to the much denser cusps observed in CDM (Navarro & Frenk, 1996) and WDM (Mocz et al., 2019).
@@
In dSph galaxies it is known that stellar motions are dominated by the dark matter gravitational potential and so we expect some visible differences for stars in these galaxies depending on the nature of dark matter. In addition, the formation and early evolution of galaxies and filaments will also be sensitive to the nature of DM, arising in particular from the suppressed form of the high-k power spectrum for the WDM/DM models. The wave behaviour and the quantum pressure of DM provide different virialized DM halo structures and prominent soliton cores, unlike the smooth cores of WDM. While for CDM and WDM cuspy profiles are expected (Navarro & Frenk, 1996; Lovell et al., 2014), DM simulations of merging DM haloes clearly show the formation of soliton structures at the de Broglie wavelength scale (Schive et al., 2014a; Mocz et al., 2017; Schwabe et al., 2016), and are prominent in terms of the core density, standing above the surrounding DM halo. This is in contrast to the early fragmentation of filaments and cuspy haloes in CDM (Mocz et al., 2017, 2019, 2020). Mocz et al. (2019, 2020) demonstrated that baryonic feedback does not have a significant effect in halos of to for redshifts , in the sense of being unable to soften the cuspy profiles of CDM/WDM to produce cores. Consequently, neither CDM/WDM are able to explain the origin of the claimed cores from the dynamical studies of dSph galaxies (Read & Gilmore, 2005; Amorisco et al., 2013).
@@
In this paper we compare the galaxy formation process between CDM, WDM, and DM, in the high-resolution cosmological simulations by Mocz et al. (2019, 2020, 2017). Here, we will consider the distribution of stars modelled in these simulations, where in the WDM and DM contexts these simulations have already reported stars to be born along dense dark matter filaments, tracing dark matter, and highlighted as “a smoking gun” signature of DM (Mocz et al., 2019). Here we will further examine these lowest redshifts, focusing on the question of core-halo structure for comparison with new deep imaging and spectroscopic data for the Milky Way and other local group dwarfs. For these objects, the data now indicates non-smooth profiles in these dark matter-dominated dwarfs as this stellar structure may potentially have important implications for the nature of dark matter. Moreover, we will look into the different evolution of the stellar populations and the impact of the three DM models on their spatial distribution.
@@
The structure of this paper is as follows: firstly, in sections 2 and 3, we describe the radial structure and internal dynamics of the DM halo for comparison with the data. Next, in section 4, we present the simulation setup. In section 5, we focus on presenting all the results of the work, with particular emphasis on the main finding related to the detection of the core-halo structure. Finally, in section 6, we discuss our conclusions regarding the non-detection of the core-halo structure for CDM and its implications.
2 The Wave Dark Matter Halo
@@
Ultralight bosons, such as Axions, have been considered as an ingredient in describing the DM (Widrow & Kaiser, 1993; Hu et al., 2000). In the absence of any self-interactions, the boson mass is the only free parameter that describes the DM. For a sufficiently light boson mass, the de-Broglie wavelength exceeds the mean free path leading the Axion to satisfy the ground state condition for a Bose-Einstein condensation. In the comoving coordinate, the governing equation for Axion reads as:
where is the wave function, is the gravitation potential and (1/(1+z)) is the cosmological scale factor. The system is normalized to the time scale , and to the scale length , with , where is the current density parameter (Widrow & Kaiser, 1993).
@@
Recently, it has proved possible with advanced GPU computing to make reliable, high dynamic range cosmological simulations that solve the above equations, (Schive et al., 2014a; Schwabe et al., 2016; Mocz et al., 2017; May & Springel, 2021). These evolve to produce large-scale structures indistinguishable from CDM, but with virialized halos characterized by a solitonic core in the ground state that naturally explains the dark matter-dominated cores of dSph galaxies (Schive et al., 2014b). Another important feature arising from the simulations is that the central soliton is surrounded by an extended halo with a “granular” texture on the de-Broglie scale, due to interference of excited states, but which when azimuthally averaged follows the Navarro-Frenk-White (NFW) density profile (Navarro & Frenk, 1996; Woo & Chiueh, 2009; Schive et al., 2014a, b).
@@
The fitting formula for the density profile of the solitonic core in a DM halo is obtained from cosmological simulations (Schive et al., 2014a, b) as:
@@
where is the boson mass and is the solitonic core radius. The latter scales with the product of the galaxy mass and boson mass, obeying the following scaling relation which has been derived from simulations (Schive et al., 2014b):
@@
Beyond the soliton, at radii larger than a transition radius (), the simulations also reveal that the halo roughly looks like NFW in shape, presumably reflecting the non-relativistic nature of condensates beyond the de Broglie scale, and therefore the total density profile can be written as:
In detail, the scale radius of the solitonic solution, which represents the ground state of the Schrodinger-Poisson equation, is related to the size of the halo through the uncertainty principle. From cosmological simulations, the latter is found to hold non-locally, relating a local property with a global one (for more details we refer to Schive et al. (2014b)).
@@
3 Dynamical Model of Galaxies in Wave Dark Matter Halo

### v2

@@
11email: alvaro.pozolarrocha@bizkaia.eu; tom.j.broadhurst@gmail.com; 22institutetext: DIPC, Basque Country UPV/EHU, E-48080 San Sebastian, Spain 33institutetext: Center for Astrophysics Harvard Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 44institutetext: Department of Astrophysical Sciences, Princeton University, 4 Ivy Lane, Princeton, NJ, 08544, USA 55institutetext: Lawrence Livermore National Laboratory, 7000 East Ave, Livermore, CA 94550, USA 66institutetext: Ikerbasque, Basque Foundation for Science, Bilbao, E-48011, Spain 77institutetext: Hong Kong University of Science and Technology, Institute for Advanced Study and Department of Physics, IAS TT WF Chao Foundation Professor, Hong Kong 88institutetext: Energetic Cosmos Laboratory, Nazarbayev University, Nursultan, Kazakhstan 99institutetext: Paris Centre for Cosmological Physics, APC, AstroParticule et Cosmologie, Université de Paris, CNRS/IN2P3, CEA/lrfu,Université Sorbonne Paris Cité, 10, rue Alice Domon et Leonie Duquet, 75205 Paris CEDEX 13, France Emeritus 1010institutetext: Physics Department LBNL, University of California at Berkeley CA 94720 Emeritus 1111institutetext: Department of Physics, Kavli Institute for Astrophysics and Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA 1212institutetext: Black Hole Initiative at Harvard University, 20 Garden Street, Cambridge, MA 02138, USA
Galaxy formation with wave/fuzzy dark matter: The core-halo structure and the solitonic imprint
@@
Dark matter-dominated cores have long been claimed for the well-studied local group dwarf galaxies. More recently, extended stellar halos have been uncovered around several of these dwarfs through deeper imaging and spectroscopy. Such core-halo structures (inner flat core and a characteristic asymptotic outer halo profile) are not a feature of conventional cold dark matter (CDM). In contrast, smooth and prominent dark matter cores are predicted for wave/fuzzy dark matter (DM). The question arises as to what extent the visible stellar profiles should reflect this dark matter core structure. Here we compare cosmological hydrodynamical simulations of CDM, “WDM” (model used as a proxy for DM) DM, aiming to predict the stellar profiles for these three DM scenarios. We show that cores surrounded by extended halos are distinguishable for DM, where the stellar density is enhanced in the core due to the presence of the relatively dense soliton. Our analysis demonstrates that, in our simulations, a distinctive core-halo structure does not appear in the case of CDM in the DM, gas, or stars. Whereas we do find a core-halo transition for DM, gas, and stars for DM, and the scale of this transition is in line with the predicted core radius set by the soliton scale anticipated for the adopted boson mass of 2.5eV. The presence of a core-halo structure in the stellar profile for Galaxy 1 for DM is visible for the most massive and the first galaxy to form in the simulation. Clearly, further simulations are needed to establish how strict this possible relationship is between the DM and stellar core-halo profile as a potential observational discriminator. Furthermore, we observe the anticipated asymmetry for DM due to the soliton’s motion (jumping and random walk), a distinctive characteristic not found in the symmetric distributions of stars in the warm and CDM models.
Key Words.:
cosmology – dark matter – galaxies1 Introduction
@@
Dark matter is understood to be nonrelativistic, i.e., cold, as required to form galaxies gravitationally and to explain the spectrum of CMB fluctuations in detail (Aghanim et al. 2018). However, the standard heavy particle interpretation for cold dark matter (CDM) faces a stringent laboratory absence of any new particle signature (Aprile et al. 2018; Sirunyan et al. 2020) and several inconsistencies have emerged between CDM predictions and the properties of dwarf galaxies (Moore 1994; de Blok 2010; Marsh & Pop 2015; Klypin et al. 1999; Safarzadeh & Loeb 2021). Warm dark matter (“WDM”) is defined to be relativistic initially with a kiloelectron volt-scale mass, the 1.4keV mass scale used here for the “WDM” simulation induces the power spectrum suppression, due to early relativistic free streaming, but becomes nonrelativistic later via the cosmological expansion, so that the “WDM” simulation is closely approximated by N-body simulations. The “WDM” simulations analyzed here were performed using a DM initial power spectrum, without including the dynamical quantum potential. This approximation reproduces the small-scale cutoff expected for a thermal relic “WDM” particle of mass 1.4 keV, corresponding to the boson mass of eV used in the DM simulations of Mocz et al. (2020). Alternatively, for very light bosons of , an inherently nonrelativistic possibility of dark matter as a Bose-Einstein condensation (Widrow & Kaiser 1993; Hu et al. 2000), originally termed fuzzy dark matter (FDM), emerged from the first simulations in this context, which predict prominent solitonic standing wave cores surrounded by pervasive interference on the de Broglie wavelength scale within all galaxies and filaments (Schive et al. 2014a; Mocz et al. 2017; Veltmaat et al. 2018; Hui 2020). This was then called wave/fuzzy dark matter (DM) (Schive et al. 2014a; Hui et al. 2021). While DM predicts large-scale structure that are almost identical to CDM in terms of the network of filaments and clusters (Schive et al. 2014a; Mocz et al. 2017), the small-scale structure of DM halos is very different in the DM paradigm, as DM cannot be confined to less than the de Broglie scale. Consequently, in DM, with a light boson mass of , the dwarf galaxy formation is suppressed and there is also a prominent solitonic core (Schive et al. 2014a, b; Mocz et al. 2017; Veltmaat et al. 2018; Niemeyer 2020) in the galaxy.
@@
Many independent claims of dark matter and stellar cores in deep images and spectroscopy are now available for the well-studied Local Group dwarf spheroidal galaxies, dSph, providing one of the main sources of tension for CDM for which the lack of any predicted cores reflects the scale-free nature of the simplest collisionless DM scenario. More recently the presence of extended stellar halos has been uncovered in several local dwarfs, independently, via deep imaging and spectroscopy (Chiti et al. 2021; Collins et al. 2021). A joint analysis of the available data now suggests that such a core-halo structure with a sharp density transition between the core and the halo is a common feature of dSph galaxies and also of the smaller new class of ultra-faint galaxies (UFGs) (Pozo et al. 2024, 2023). This core-halo behavior (Chan et al. 2020a) motivates the examination of hydrodynamical simulations with a comparison of the differing DM contenders as a possible distinguishing feature.
@@
For cold dark matter (CDM), the scale-free formation of DM structures is predicted to extend to arbitrarily small scales, reaching lower halo masses with higher DM concentrations. In contrast, small-scale suppression of the power spectrum is inherent to both DM and “WDM”. In the case of DM, this suppression limits the minimum scale of structure to above the de Broglie length, which is determined by the boson mass. Consequently, DM structures below are suppressed for a choice of boson masses of eV, motivated by the observed kpc scale of dwarf galaxy cores. This limitation has a direct consequence for the formation timescales between CDM and DM, and results in a delay in galaxy formation for DM compared to CDM. In both cases, we observe that the formation of the first filamentary structures occur earlier in CDM, where they are comprised of low-mass subhalos. In contrast, in DM/“WDM” there is no fragmentation along the filaments, due to the small-scale cutoff of the power spectrum. Instead, dark matter is distributed more continuously along these filaments (May & Springel 2023; Dome et al. 2023).
@@
The main difference between “WDM” and DM is that “WDM” exhibits a clear caustic structure in the distribution of dark matter, while DM is characterized by wave interference patterns that manifest as fringes in simulations. These fringes are now understood to generate turbulent density structures within halos, resulting from the superposition of numerous plane waves encoding the velocity dispersion in the halo (Mocz et al. 2020). In DM these filaments persist for longer periods as the gravitational attraction from the early formation of the first halos in cold dark matter (CDM) tends to disrupt them earlier. Moreover, the prolonged lifetimes of these filaments in wave/fuzzy dark matter favor the formation of a greater number of stars within them, leading to significant differences in the location and extent of stellar profiles in CDM and DM galaxies. This also causes baryonic objects to appear more diffuse or smoothed compared to CDM (Mocz et al. 2019, 2020), indicating that they could serve as excellent tracers (as visible objects that could be used to observe the underlying DM structures) of dark matter in DM. On the scale of solitons, the quantum pressure in DM can become sufficiently strong to counteract the self-gravity of the dark matter. This results in the formation of prominent solitonic standing waves at the base of each virialized potential, characterized by a spherical soliton core of a few kiloparsecs at the center of every dark matter halo (Schive et al. 2014a), in contrast to the much denser cusps observed in CDM (Navarro & Frenk 1996).
It is remarkable how well the soliton profile of DM aligns with the empirically well-established Burkert profile, as demonstrated in the extensively studied Fornax dwarf spheroidal galaxy Schive et al. (2014a). However, the scaling of the core radius with galaxy mass reported by Burkert (2020) from profile fitting to low- and intermediate-mass galaxies exhibits a slope that is opposite in sign to that predicted by DM, as highlighted by Burkert (2020). Moreover, the relatively large cores inferred for intermediate-mass galaxies are based on HI gas dynamics. Noncircular gas motions have been recognized as a significant contribution to underestimated rotation curves (Oman et al. 2019; Downing & Oman 2023), together with resolution limitation for HI (Swaters et al. 2000) and the finite thickness of HI disks that all act to bias downward HI-based rotation speed measurements, as demonstrated using CDM-based hydro simulations with spurious large cores (Roper et al. 2023). It is important to examine the claimed scaling of galaxy mass and core radius with independent stellar-based dynamics, though this remains challenging due to the typically low stellar surface brightnesses.
@@
In dSph galaxies it is known that stellar motions are dominated by the dark matter gravitational potential, and so we expect some visible differences for stars in these galaxies depending on the nature of dark matter. In addition, the formation and early evolution of galaxies and filaments is also sensitive to the nature of DM, arising in particular from the suppressed form of the high-k power spectrum for the “WDM”–DM models. The wave behavior and the quantum pressure of DM provide different virialized DM halo structures and prominent solitonic cores, unlike the CDM–WDM cuspy profiles (without inner cores) expectations(Navarro & Frenk 1996) (even smooth cores of “ WDM are claimed (Lovell et al. 2014; Macciò et al. 2012; Bode et al. 2001)), DM simulations of merging DM halos clearly show the formation of soliton structures at the de Broglie wavelength scale (Schive et al. 2014a; Mocz et al. 2017; Schwabe et al. 2016), and are prominent in terms of the core density, lying above the surrounding DM halo. This is in contrast to the early fragmentation of filaments and cuspy halos in CDM (Mocz et al. 2017, 2019, 2020). Mocz et al. (2019, 2020) postulated that baryonic feedback may not have a significant effect in halos of to for redshifts , in the sense of being unable to soften the cuspy profiles of CDM to produce cores. Consequently, CDM should not be able to explain the origin of the claimed cores from the dynamical studies of dSph galaxies (Read & Gilmore 2005; Amorisco et al. 2013).
@@
In this paper we compare the galaxy formation process between CDM, “WDM”, and DM, in the high-resolution cosmological simulations by Mocz et al. (2019, 2020, 2017). Here we consider the distribution of stars modeled in these simulations, where in the “WDM” and DM contexts these simulations have already reported stars to be born along dense dark matter filaments, tracing dark matter, and highlighted as “a smoking gun” signature of DM/“WDM” (Mocz et al. 2019). Here we further examine these lowest redshifts, focusing on the question of core-halo structure for comparison with new deep imaging and spectroscopic data for the Milky Way and other local group dwarfs.
@@
The structure of this paper is as follows. In sections 2 and 3 we describe the radial structure and internal dynamics of the DM halo for comparison with the data. Next, in section 4 we present the simulation setup. In section 5 we focus on presenting all the results of the work, with particular emphasis on the main finding related to the detection of the core-halo structure. Finally, in section 6 we discuss our conclusions regarding the nondetection of the core-halo structure for CDM and its implications. The boson masses used in this work are in strong tension with current constraints in the literature (for a discussion of these constraints, see section 6).
2 The wave/fuzzy dark matter halo
@@
Ultralight bosons, such as ALPs (axion-like particles), have been considered as an ingredient in describing the DM (Widrow & Kaiser 1993; Hu et al. 2000). In the absence of any self-interactions, the boson mass is the only free parameter that describes the DM. For a sufficiently light boson mass, the de-Broglie wavelength exceeds the mean free path, which leads the ALP to satisfy the ground state condition for a Bose-Einstein condensation. The governing equation for DM is the Schrödinger-Poisson (SP) equation, which in comoving coordinates reads as:
where is the wave function (see Schive et al. (2014a) for the explanation of the interpretation of the wave function, originally introduced by Widrow & Kaiser (1993), to describe the mean field behavior of the wave function in this context), is the gravitation potential and =(1/(1+z)) is the cosmological scale factor, where z is the redshift.The comoving length is normalized to ( , the time normalized to and the wave function normalized to , where is the present Hubble parameter, is the matter density parameter and the background mass density, (Schive et al. 2014b). Finally, the physical interpretation of the wave function in DM is that the density is given by .
@@
Recently, it has become possible to perform high dynamic range cosmological simulations that solve the above equations (Schive et al. 2014a; Schwabe et al. 2016; Mocz et al. 2017; May & Springel 2021), where GPU computing has allowed an efficient adaptive mesh refinement (Schive et al. 2014a). These evolve to produce large-scale structures indistinguishable from CDM, but with virialized halos characterized by a solitonic core in the ground state that naturally explains the dark matter-dominated cores of dSph galaxies, previously analyzed in many dSph and UFD of the Local Group (Pozo et al. 2024, 2023, 2021, 2022; Schive et al. 2014b) . Another important feature arising from the simulations is that the central soliton is surrounded by an extended halo with a granular texture on the de-Broglie scale, due to interference of excited states, but which when azimuthally averaged follows the Navarro-Frenk-White (NFW) density profile (Navarro & Frenk 1996; Woo & Chiueh 2009; Schive et al. 2014a, b).
@@
The fitting formula for the density profile of the solitonic core in a DM halo is obtained from cosmological simulations (Schive et al. 2014a, b) as
@@
where is the boson mass and is the solitonic core radius. The latter scales with both the halo mass and the boson mass, obeying the following scaling relation which was derived from the simulations (Eq. 7 (Schive et al. 2014b))
@@
where is the halo mass and . is defined as the matter density parameter as a function of redshift, normalized to the critical density at that time. Beyond the soliton, at radii larger than a transition radius (), the simulations also reveal that the halo roughly resembles NFW in shape, presumably reflecting the nonrelativistic nature of condensates beyond the de Broglie scale, and therefore the total density profile can be written as
where is chosen such that the inner solitonic profile matches the outer NFW-like profile at approximately , and is the scale radius. In detail, the scale radius of the solitonic solution, which represents the ground state of the Schrödinger-Poisson equation, is related to the size of the halo through the uncertainty principle.
@@
3 Dynamical model of galaxies in wave/fuzzy dark matter halo